OpenStudy (anonymous):

A sphere and a cube have equal surface areas. The ratio of the volume of the sphere to that of the cube isa. √π : √6b. √π : √6c. √π : √3d. √6 : √π

OpenStudy (anonymous):

The marked price of a table is Rs. 3000 and is available at successive discounts of 20% and 10% respectively. If there is an additional discount of 5% on cash payment, then what is the cash price of the table?a. Rs. 2400b. Rs. 2160c. Rs. 2100d. Rs. 2052

OpenStudy (anonymous):

Price of food grains have risen by 10% and of other items of consumption by 15%. If the ratio of an employee's expenditure on food grains and other items is 2:5, by how much should his salary be increased so that he may maintain the same level of consumption as before, assuming that his present salary is Rs. 3500?a. Rs. 300b. Rs. 350c. Rs. 375d. Rs. 475

OpenStudy (anonymous):

please solve this paper

OpenStudy (jtvatsim):

try to solve some of the questions on your own, we'll be more than happy to check or explain concepts that confuse you...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

please reply me

OpenStudy (anonymous):

dear do u know the formula of vuliom and surface ?

OpenStudy (anonymous):

volume *

OpenStudy (anonymous):

sphere : volume :\[r^{3} \pi \frac{ 4 }{ 3 }\] surface: \[ 4\pi r^{ 2} \]

OpenStudy (anonymous):

cube : sufrace 6a^2 volume a^3

OpenStudy (anonymous):

use these clue to solve first question u need two equation to solve this problem ur first clue is the surface of both cube and sphere are the same so u can find a relationship between r and a and use this relationship in the next part .

OpenStudy (anonymous):

thanks for medal , i'm here for next 30 min . u can ask any quastion

OpenStudy (anonymous):

the ratio 435:25 is equal to 4:1 how?

OpenStudy (anonymous):

sri i didn't get it . what u mean by the ratio 435/25 is 4/1 that's not correct

OpenStudy (anonymous):

actually u can't divide 435 over 4 this has remine

OpenStudy (anonymous):

see this attachment

OpenStudy (anonymous):

which page ? which qustion?

OpenStudy (anonymous):

i think its in 4th page

OpenStudy (anonymous):

19th question

OpenStudy (anonymous):

u guys really great.hats off to open study

OpenStudy (anonymous):

dear this qustion is totally wrong . 435 is 87 times of 5 and 25 is 5 times of 5 u can cansel 5 in nom and dem and u reach 87/5

OpenStudy (anonymous):

\[\frac{ 87*5 }{ 5*5 }\frac{ 435 }{ 25 }=\]dear this qustion is totally wrong . 435 is 87 times of 5 and 25 is 5 times of 5 u can cansel 5 in nom and dem and u reach 87/5

OpenStudy (anonymous):

the equal symbol should be in the middle

OpenStudy (anonymous):

thank u very much

OpenStudy (anonymous):

ur welcome ,pleasure

OpenStudy (anonymous):

The marked price of a table is Rs. 3000 and is available at successive discounts of 20% and 10% respectively. If there is an additional discount of 5% on cash payment, then what is the cash price of the table?a. Rs. 2400b. Rs. 2160c. Rs. 2100d. Rs. 2052

OpenStudy (anonymous):

sorry i got the answer to this question

OpenStudy (anonymous):

Price of food grains have risen by 10% and of other items of consumption by 15%. If the ratio of an employee's expenditure on food grains and other items is 2:5, by how much should his salary be increased so that he may maintain the same level of consumption as before, assuming that his present salary is Rs. 3500?a. Rs. 300b. Rs. 350c. Rs. 375d. Rs. 475

OpenStudy (anonymous):

Price of food grains have risen by 10% and of other items of consumption by 15%. If the ratio of an employee's expenditure on food grains and other items is 2:5, by how much should his salary be increased so that he may maintain the same level of consumption as before, assuming that his present salary is Rs. 3500?a. Rs. 300b. Rs. 350c. Rs. 375d. Rs. 475